OpenStudy (anonymous):

probability question: choose 19 integers \[ 0 le X_r le 10 \] find the probability their sum is greater than 120

OpenStudy (anonymous):

\[0 \le X_r \le 10\]

OpenStudy (anonymous):

i have an answer but i got it through a fairly brute force method

OpenStudy (anonymous):

i would be really interested if there was a nice/general solution

OpenStudy (anonymous):

hmm... maybe to find out the formula would be possible to think like that: First consider groups of 19 integers, all equal that do summ more than 120: 1010101010.... 9999... 88888... 77777.... Later for each of this group there would be a margin that the have over 120: 19*10-120=70 19*9-120=51 .... this margin could be summed by numbers smaller , but all equal, then the original number. This way you also will get groups of equal numbers with new margins. I think this way you can manage to get all possible combinations and work out recursive type formula for the general solution.

OpenStudy (anonymous):

this is the manner in which i proceeded, which might also be a possible foot-in to a formula (notice pascal triangle diagonals in rows) each column is an extra integer added, each row shows the probability of the sum being 1,2,3,4,5,... etc

OpenStudy (anonymous):

looks promising, :)

OpenStudy (anonymous):

Is the answer around 3.8%?

OpenStudy (anonymous):

0.031975056

OpenStudy (anonymous):

this was the original problem this came from: Consider a computer game where you play a level 1 character with 150 HP. Each time your character levels, you have a random chance of gaining between 20 to 30 HP (So you can be level 2 with a HP anywhere between 170 to 180. At level 20: given that HP is discrete (takes integer values) a) Calculate the range of HP your character could have b) Calculate the probability your character has over 650 HP